What is the work done by GRAVITY if a 50kg block is pushed along a 4m horizontal track?

1
Expert's answer
2022-02-24T09:05:07-0500

The work done by the force of gravity

"W=Fd\cos\theta=50*9.8*4\cos 90^\circ=0"